-3x-18y=6
6x-y=14

Answer by stanbon(75887)

-3x-18y=6
6x-y=14
------
Modify for elimination:
Multiply thru the top equation by 2:
-----
-6x - 36y = 12
6x - y = 14
-------------------
Add to get:
-37y = 26
---
y = -26/37
----------------------------
Substitute for "y" and solve for "x" in the bottom equation:
6x + (26/37) = 14
------
222x + 26 = 518
222x = 492
x = 82/37
================
Cheers,
Stan H.
=================